OpenStudy (lovetostudy):

The math club raised money for its spring banquet by washing vehicles. The club charged $3 per car and $5 per truck. The club earned a total of total of $ 550 for washing 122 cars and trucks combined. How many cars did the club wash?

OpenStudy (anonymous):

So, you have to form a system of equations

OpenStudy (johnweldon1993):

this is a system of equations 3 dollars per car means 3c (c being cars washed) and 5 dollars per truck washed means 5t (t being trucks washed) and that total is 550....so 3c + 5t = 550 you know they washed 122 cars (c) and trucks (t) combined...so c + t = 122 so you have 3x + 5t = 550 c + t = 122 do you know how to solve a system of equations?

OpenStudy (johnweldon1993):

sorry should be 3c in that first equation!!***

OpenStudy (lovetostudy):

I mean, I am pretty sure my Algebra teacher taught this in class. But, I just can never remember or comprehend what she teaches. Mathematics is my total weakness :(

OpenStudy (anonymous):

good explaining johnweldon :)

OpenStudy (johnweldon1993):

thank you @kelliegirl33 :) and okay @lovetostudy we'll walk through this step by step :)

OpenStudy (johnweldon1993):

so we have 3c + 5t = 550 c + t = 122 what we want to do.....is solve for the number of CARS washed correct? this means....that you eventually want to solve for "c" take c + t = 122 solve for t here to isolate t (get by itself)....you would subtract "c" from both sides...... and get t = 122 - c right? with me so far?

OpenStudy (lovetostudy):

yes.

OpenStudy (lovetostudy):

we're solving for C (cars) by isolating T (trucks) , correct?

OpenStudy (johnweldon1993):

CORRECT! so not...you take that t.....and what it equals (122 - c)...and plug it into that first equation 3c + 5t = 550 and plug in (122-c) wherever you see a "t" so 3c + 5(122 - c) = 550 still with me?

OpenStudy (johnweldon1993):

***notice that the ONLY varible left.....is C...which IS what we're solving for...that is exactly why we isolate T in the first part :)

OpenStudy (lovetostudy):

Would we need to use distributive property next?

OpenStudy (johnweldon1993):

yes indeed we would :) so what does our equation look like after that?

OpenStudy (lovetostudy):

3c+610=550?

OpenStudy (anonymous):

you forgot the 5 x -c

OpenStudy (johnweldon1993):

close....what is the distributive property? where you have 5(122 - c) you would multiply the 5 by 122....AND THEN 5 times -c as well right?

OpenStudy (lovetostudy):

I know I have to do something with the -c that's in the parentheses, but I forgot :(

OpenStudy (lovetostudy):

Yes.

OpenStudy (johnweldon1993):

so you do exactly that 5 times 122....is the 610 5 times -c is?

OpenStudy (lovetostudy):

-5c

OpenStudy (johnweldon1993):

perfect! so altogether we have 3c + 610 -5c = 550 where would we go from here?

OpenStudy (lovetostudy):

Um, like terms? :o

OpenStudy (johnweldon1993):

yes! ....meaning?

OpenStudy (lovetostudy):

we add 3c + -5c which would equal ... -2c?

OpenStudy (johnweldon1993):

exactly! so -2c + 610 = 550 what now?

OpenStudy (lovetostudy):

We subtract 610 from itself so it cancels out and then subtract that from 550?

OpenStudy (johnweldon1993):

mmhmmm...whick leaves us with?

OpenStudy (johnweldon1993):

which*

OpenStudy (lovetostudy):

-2c+610=550 -610 -610 then, -2c -60 ----- = ------ -2c -2c the answer is ... 30? :D

OpenStudy (johnweldon1993):

very good! :) but for future reference....we dont divide -2c and -60 by -2"c" we just divide it by -2 if we divide them by -2c we'll have -2c = -60 ---------- -2c -2c 1 = -60/-2c not the same just for the future....but very good :)

OpenStudy (lovetostudy):

Ohh, ahha. Sorry, Typo. I was doing it on a loose sheet of paper and mistakenly included the variable. Oops.

OpenStudy (johnweldon1993):

not a problem ..just wanted to make sure! :) very good work @lovetostudy :)

OpenStudy (lovetostudy):

Thank you :) Just one more question, (sorry to bother) If my final answer is 30, then why does it not appear as an answer choice on my worksheet? :o That's very odd.

OpenStudy (johnweldon1993):

what are your answer choices?

OpenStudy (lovetostudy):

1) 36 2) 50 3) 72 4) 86

OpenStudy (johnweldon1993):

hang on let me double check

OpenStudy (johnweldon1993):

ummm...no it should definitely be 30 cars....which corresponds to 92 trucks...and to check it 3(30) + 5(92) = 550 90 + 460 = 550 550 = 550 its correct...so your answer choices must be wrong

OpenStudy (johnweldon1993):

let me get a different opinion first though

OpenStudy (johnweldon1993):

@jim_thompson5910 mind looking this over for me?

OpenStudy (lovetostudy):

Okay :)

jimthompson5910 (jim_thompson5910):

I'm getting c = 30 and t = 92 as well so there has to be a typo somewhere

OpenStudy (johnweldon1993):

thank you very much @jim_thompson5910 :)

OpenStudy (lovetostudy):

Thank you, Jim

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (lovetostudy):

I'm guessing there IS a typo :o Thank you so so much John. You've really guided me to see the light at the end of the dark tunnel. Haha Greatly appreciate your guidance. :)

OpenStudy (lovetostudy):

Oh!! It was my mistake all along . The total earned was 510 instead of 550 -.-

jimthompson5910 (jim_thompson5910):

ok so one of those answer choices is correct assuming the 550 changes to 510

OpenStudy (lovetostudy):

Yes.

jimthompson5910 (jim_thompson5910):

what do you get for c this time

OpenStudy (johnweldon1993):

right do this again the same way...but with 510

OpenStudy (lovetostudy):

It's 50. I found the answer thanks to your help & refreshing my mind of how to solve these types of problems :D

OpenStudy (johnweldon1993):

great job :)

OpenStudy (lovetostudy):

Oh my goodness, I am so proud of myself. I would've never done it without you :) and Jim :)

jimthompson5910 (jim_thompson5910):

lol didn't do anything really, but yw

OpenStudy (johnweldon1993):

lol sometimes a second opinion / confirmation is all the help you need Jim!

jimthompson5910 (jim_thompson5910):

hmm true, but you did most of the work

OpenStudy (lovetostudy):

So true.